### v3

# Redefining non-IID Data in Federated Learning for Computer Vision Tasks: Migrating from Labels to Embeddings for Task-Specific Data Distributions

Abstract
Federated Learning (FL) represents a paradigm shift in distributed machine learning (ML), enabling clients to train models collaboratively while keeping their raw data private. This paradigm shift from traditional centralized ML introduces challenges due to the non-iid (non-independent and identically distributed) nature of data across clients, significantly impacting FL’s performance. Existing literature, predominantly model data heterogeneity by imposing label distribution skew across clients. In this paper, we show that label distribution skew fails to fully capture the real-world data heterogeneity among clients in computer vision tasks beyond classification. Subsequently, we demonstrate that current approaches overestimate FL’s performance by relying on label/class distribution skew, exposing an overlooked gap in the literature. By utilizing pre-trained deep neural networks to extract task-specific data embeddings, we define task-specific data heterogeneity through the lens of each vision task and introduce a new level of data heterogeneity called embedding-based data heterogeneity. Our methodology involves clustering data points based on embeddings and distributing them among clients using the Dirichlet distribution. Through extensive experiments, we evaluate the performance of different FL methods under our revamped notion of data heterogeneity, introducing new benchmark performance measures to the literature. We further unveil a series of open research directions that can be pursued.
Index Terms:
Federated learning, computer vision, data heterogeneity, label distribution skew, non-iid data.
I Introduction
How does data heterogeneity impact the performance of federated learning (FL) in computer vision tasks? Are current approaches adequately capturing the effects of data heterogeneity on FL performance? Surprisingly, our short answer is No. In this paper, we reveal a notable oversight in the literature – namely, that the performance of existing FL strategies in computer vision has been overestimated in the presence of data heterogeneity among the participating clients.
I-A Overview and Literature Review
FL has emerged as a promising paradigm for distributed machine learning (ML), which enables multiple clients to collaboratively train ML models without sharing raw data, thus preserving user privacy[1, 2, 3]. This approach offers an alternative to traditional centralized ML techniques, where data must be transferred to a central server, often resulting in significant communication overhead and privacy concerns. The training process in FL takes place over several global rounds. In each round, clients first train ML models locally on their datasets, typically using several iterations of stochastic gradient descent (SGD). They then send the updated local models to a central server, which aggregates the models (usually through averaging) to form a new global model. This global model is subsequently broadcast back to the clients to start the next round of local training. Implementing FL in practice faces unique challenges, largely due to the inherent non-iid (non-independent and identically distributed) or heterogeneous nature of data across clients [2, 3, 4, 5, 6, 7, 8]. This data heterogeneity has a significant impact on FL performance. Specifically, as data non-iid-ness (or heterogeneity) increases, local models of clients become biased toward their local datasets, negatively affecting the convergence of the global model [3].
Effective evaluation of the robustness of FL algorithms thus necessitates examining how data heterogeneity influences FL performance. To this end, state-of-the-art methods in FL that concern modeling the data heterogeneity have mainly emulated data heterogeneity across clients through imposing label distribution skew [^1] The label distribution skew approach involves partitioning data in such a way that clients receive data with imbalanced label distributions [1, 9]., largely applied to computer vision classification tasks [2, 3, 10, 11, 7, 12, 9, 13, 5]. In particular, when FL first emerged as a decentralized ML paradigm, researchers emulated data heterogeneity through distributing data points based on their labels across clients to impose label distribution skew. This approach became a widely used method to simulate data heterogeneity in subsequent FL research [2, 10, 7, 12]. For example, in a handwritten digits classification task (MNIST dataset [14]), one client might have a local dataset mostly consisting of the digits “0” and “1”, while another client might only have images of the digits “7” and “8” [2]. As the number of labels per clients decreases, each client’s dataset becomes less representative of the global dataset, emulating a more non-iid data distribution. While this fundamental approach introduces some degree of data heterogeneity, there was a pressing need for mathematical and unified approaches to introduce and measure the level of data heterogeneity in FL. A major step in this direction was taken by utilizing the Dirichlet distribution to impose label distribution skew among clients in FL [3, 12, 11, 9, 13, 5, 7], enabling gradual heterogeneity in label/class distribution across the clients by fine-tuning Dirichlet concentration parameter.
Nevertheless, the above-mentioned class-based (equivalently, label-based) data heterogeneity may not be a prefect match for other types of tasks, specifically computer vision tasks beyond classification tasks, such as surface normal estimation, depth estimation, edge detection, and object detection [15, 12] (see Fig. 2 for a list of these tasks). Intuitively, the ineffectiveness of class-based data heterogeneity methods for generic computer vision tasks stems from the fact that non-classification vision tasks do not necessarily perceive data distribution in the same way as classification tasks. In fact, their perception of data heterogeneity can be label-independent and majorly different from each other: an aspect largely overlooked in the existing FL literature. For instance, the “2D Edges” task may be more sensitive to the configuration of edges within an image rather than the image’s label. In this context, two datapoints with the same label (e.g., both with “Living Room” label) may exhibit highly different edge structures, rendering them heterogeneous from the perspective of the “2D Edges” task. Conversely, two datapoints with different labels (e.g., “Living Room” and “Kitchen”) might have very similar edge structures, making them homogeneous for this specific task despite having different labels. This highlights the need for methodologies in FL that emulate data heterogeneity based on the unique perspectives of vision tasks to data.
Building on the above intuition, in this paper, we propose the following fundamental hypothesis:
As we demonstrate in this paper, the performance of FL for generic computer vision tasks has been overestimated in prior studies [16, 17, 18, 19]. These studies largely depend on class-based heterogeneity, which fails to accurately reflect the data distribution in real-world clients’ local datasets. Consequently, the actual performance and robustness of FL, and its variants, in handling data heterogeneity for generic computer vision tasks remain largely unexplored, highlighting a significant gap in the current literature.
To bridge this gap, we redefine the notion of heterogeneity in FL for computer vision tasks as follows:
Instead of inducing heterogeneity solely from a raw data perspective, we propose a migration towards assessing data heterogeneity through the unique perspective of each task to the data after applying task-specific transformations on data.
To achieve the above objective, for the first time in the literature, we introduce a new layer of data heterogeneity, which we call embedding-based data heterogeneity. To model embedding-based data heterogeneity (i.e., embedding distribution skew), we leverage a pre-trained deep neural network (DNN) for each task and feed forward data points through this network. We then exploit the fact that the initial layers of the DNN of each task extract low-level features. As the data progresses through deeper layers, the network extracts higher-level representations that are increasingly specific to the vision task of interest. As we reach the layer just prior to the last layer (i.e, penultimate layer), embeddings that encapsulate a comprehensive view of the data from the perspective of the task are obtained, integrating both low-level and high-level features. We then leverage K-means to cluster datapoints based on these embeddings. Subsequently, we utilize a Dirichlet distribution to assign data points of each cluster to different clients to emulate task-specific data heterogeneity.
I-B Summary of Contributions
Our main contributions can be summarized as follows:
$\bullet$
Through comprehensive experimental studies, we demonstrate that prior methods that impose label distribution skew to emulate data heterogeneity across clients, largely applied to computer vision classification tasks, have led to an overestimation of the performance of FL for generic computer vision tasks other than classification tasks.
$\bullet$
We demonstrate that the concept of data heterogeneity can vary significantly across different tasks. This calls for a migration from one-size-fits-all approach in the literature of FL where data heterogeneity is always induced via imposing label distribution skew across clients datasets.
$\bullet$
To uncover the actual performance of FL for computer vision tasks under data heterogeneity, for the first time, we introduce task-specific data heterogeneity, which we refer to as embedding-based data heterogeneity. By revisiting FL with our refined concept of data heterogeneity, we reveal that FL’s performance can be significantly impacted by task-specific data heterogeneity for a variety of vision tasks.
$\bullet$
We highlight the significance of task similarity in FL, illustrating that similar tasks tend to perceive data heterogeneity in a consistent manner, while differing tasks may interpret the same data distribution across clients with varying degrees of heterogeneity. Through this nuanced observation we unveil a series of novel research directions on addressing data heterogeneity for vision tasks in single-task and multi-task FL settings.
Fig. 1: Dirichlet-generated data distribution for $N=20$ clients and $10$ classes with homogeneous prior probabilities (i.e., $p_{i}=1/|\mathcal{C}|,\leavevmode\nobreak\ \forall 1\leq i\leq|\mathcal{C}|$). In the case of $\alpha=10^{6}$, the data is homogeneously distributed across clients, whereas in the case of $\alpha=0.1$, the majority of each client’s data is consisted of two labels at most. The decrease in the homogeneity of the clients’ data is also observed by comparing the data distributions generated by $\alpha=10^{3}$ and $\alpha=10$.
II Preliminaries and Motivations
In this section, we first model an FL system in Sec. II-A, describing the operations performed at both the client and server levels. We then discuss the current approaches for modeling data heterogeneity in FL in Sec. II-B, where we propose an experiment to evaluate these methods for computer vision tasks and highlight their limitations.
II-A FL System Modeling
We consider executing FL over a set of clients denoted by $\mathcal{N}$. Each client $n\in\mathcal{N}$ is assumed to possess a local dataset $\mathcal{D}_{n}$ with size $D_{n}=|\mathcal{D}_{n}|$. We also let $\mathcal{D}=\bigcup_{n\in\mathcal{N}}\mathcal{D}_{n}$ with size $D=|\mathcal{D}|=\sum_{n\in\mathcal{N}}D_{n}$ denote the global dataset, encompassing the datasets distributed across the clients. For an ML task of interest, we let $\ell(\bm{\omega};d)$ denote the loss of the task under arbitrary ML model parameter $\bm{\omega}$ and datapoint $d$. Subsequently, the local loss function of each client $n$ is given by $\mathcal{L}_{n}(\bm{\omega})=\frac{1}{D_{n}}\sum_{d\in\mathcal{D}_{n}}\ell(\bm{\omega};d)$. The goal of FL is to find the optimal model parameter $\bm{\omega}^{\star}$ to minimize the global loss function $\mathcal{L}(\bm{\omega})=\frac{1}{D}\sum_{d\in\mathcal{D}}\ell(\bm{\omega};d)$ for the ML task of interest, which can be mathematically represented as
$\bm{\omega}^{\star}=\operatorname*{arg\,min}_{\bm{\omega}}\mathcal{L}(\bm{\omega})\equiv\operatorname*{arg\,min}_{\bm{\omega}}\sum_{n\in\mathcal{N}}\frac{D_{n}\mathcal{L}_{n}(\bm{\omega})}{D}.$ (1)
To achieve $\bm{\omega}^{\star}$, FL follows an iterative approach consisting of a series of local model training rounds at the clients and global model aggregation rounds at the server. During each FL global aggregation round $t\in\{0,\cdots,T-1\}$, the central server first broadcasts the global model, $\bm{\omega}^{(t)}$, to all clients (with $\bm{\omega}^{(0)}$ typically initialized randomly). Each client $n$ will then initialize its local model with the global model as $\bm{\omega}_{n}^{(t),0}\leavevmode\nobreak\ =\leavevmode\nobreak\ \bm{\omega}^{(t)}$ and performs $K$ rounds of local ML model training via mini-batch stochastic gradient decent (SGD), each indexed by $k$, where $1\leq k\leq K$. In particular, the local model of client $n$ after the $k$-th SGD iteration during the global aggregation round $t$ is given by:
$\bm{\omega}_{n}^{(t),k}=\bm{\omega}_{n}^{(t),k-1}-\eta^{(t),k}_{n}\widetilde{\nabla}\mathcal{L}_{n}(\bm{\omega}_{n}^{(t),k-1}),$ (2)
where $\eta^{(t),k}_{n}$ is the instantaneous learning-rate. Also, $\widetilde{\nabla}$ denotes the stochastic gradient approximation defined as ~$\widetilde{\nabla}\mathcal{L}_{n}(\bm{\omega}_{n}^{(t),k-1})=\frac{1}{|\mathcal{B}_{n}^{(t),k-1}|}\sum_{d\in\mathcal{B}_{n}^{(t),k-1}}\nabla\ell(\bm{\omega}_{n}^{(t),k-1};d)$, where $\mathcal{B}_{n}^{(t),k-1}\subseteq\mathcal{D}_{n}$ represents the set of randomly sampled datapoints contained in the mini-batch.
Fig. 2: Visualization of the tasks from the Taskonomy dataset that have been used in our experiments.
After the conclusion of local model training, clients send their latest trained local ML models (i.e., $\bm{\omega}_{n}^{(t),K}$ for each client $n$) to the server, where the received models are aggregated to create a new global model as follows:
$\bm{\omega}^{(t+1)}=\sum_{n\in\mathcal{N}}\frac{D_{n}\bm{\omega}_{n}^{(t),K}}{D}.$ (3)
This new global model is then broadcast back to the clients to start the next local model training round and this iterative procedure is repeated until a desired convergence criterion is achieved. Note that the procedure formulated above describes the vanilla FedAvg method [2]. Other variations of FL, which are studied later in Sec. IV, often involve modifications to the above-described local training and global aggregation processes.
II-B Analyzing Current Approaches on Modeling Heterogeneity
II-B1 Importance and Impact of Data Heterogeneity in FL
In FL, data is collected and stored locally by the clients, never being transferred over the network. Consequently, the server has often no control over the composition or distribution of the clients’ local datasets. This leads to data heterogeneity (commonly referred to as non-iid data222“non-iid” stands for “non-independent and identically distributed.”) across clients, a fundamental challenge in FL. Specifically, when the datasets $\mathcal{D}_{n}$ and $\mathcal{D}_{n^{\prime}}$ differ significantly across clients ($n\neq n^{\prime}$), variations in local data distributions introduce local model bias during the local SGD iterations at the clients. This bias can degrade the overall performance of the FL global model. In particular, non-iid datasets result in local models that capture client-specific characteristics, which creates inconsistencies during model aggregation and exacerbates issues such as divergence and slow convergence. Modeling and addressing data heterogeneity are therefore critical for advancing FL. In this work, we aim to introduce a novel modeling approach for non-iid data in FL, leaving addressing it to future research. Specifically, we focus on highlighting the inefficiencies of existing methods for modeling non-iid data in FL, particularly for computer vision tasks. Our proposed method provides a generic framework for data heterogeneity that not only extends to computer vision tasks but also encompasses existing methods as special cases.
II-B2 State-of-the-art (SoTA) Methods in Modeling Data Heterogeneity
SoTA methods in FL that concern modeling the data heterogeneity [10, 11, 7, 12] have heavily relied on classification tasks. In this setting, the global dataset often has a limited set of class labels, which streamlines the emulation of non-iid data across the clients through imposing imbalanced class label distribution across their local datasets. More trivial approaches [9, 20] include ensembling datasets to create a heterogeneous data gathering setting [20] and partially distributing labels between clients to emulate non-iid data (i.e., each client’s local dataset contains data with a predetermined ratio from each class label)[9]. However, the most common approach used in the FL literature is a theoretically-oriented method, which utilizes the Dirichlet distribution [12, 11, 9, 13, 5, 7] to emulate non-iid data. This SoTA method considers the ratio of each available class label in each client’s local dataset as a random variable drawn from Dirichlet distribution which can be tuned with specific parameters. In the following, we first further explain this SoTA method, which induces class-based data heterogeneity across the clients in FL, and then study its effectiveness in more generic non-classification computer vision tasks.
II-B3 Creating non-iid Data via Dirichlet Distribution
In order to model the class-based data heterogeneity using the Dirichlet distribution in FL, a set of classification labels $\mathcal{C}$ with each member $c\in\mathcal{C}$ representing a class label in the global dataset is considered [^3] For example, each digit in MNIST [14] dataset.. Then, parameter $\mathbf{p}=\{p_{c}\}_{c\in\mathcal{C}}$ is used for the set of ratios of each class $c$ in the global dataset $\mathcal{D}$ which is being dispersed across clients. Afterwards, a label distribution vector is created for each client $n$’s local dataset $\mathbf{q_{n}}=\mathsf{Dir}(\alpha\mathbf{p})=\{q_{n,c}\}_{c\in\mathcal{C}}$ where $q_{n,c}$ denotes the ratio of client $n$’s dataset containing data from class label $c$. Here, $\alpha\geq 0$ is the concentration parameter which determines how closely each $\mathbf{q_{n}}$ will resemble $\mathbf{p}$, with higher $\alpha$ leading to a more similar $\mathbf{q_{n}}$ to $\mathbf{p}$. We can therefore interpret decreasing in the concentration parameter $\alpha$ as increasing the data heterogeneity across the clients. This phenomenon is visualized in Fig. 1 for an FL system with $20$ clients using the MNIST dataset [14], which contains $10$ labels (the same results apply to other datasets with $10$ labels, such as Fashion MNIST [21] and CIFAR-10 [22]). In the figure, data points with the same class label are shown in similar color tones. As illustrated, reducing $\alpha$ from $10^{6}$ to $0.1$ corresponds to increasing levels of data heterogeneity, with $\alpha=0.1$ representing an extreme case of non-iid data where most clients’ datasets contain fewer than two distinct labels. As a result, it can be construed that smaller values of $\alpha$ represent emulation of real-world scenarios where local datasets of clients induce severe local model bias during the local training rounds and negatively affect the convergence of FL global model.
Due to its fundamental reliance on class label distribution, emulation of non-iid data using Dirichlet distribution is a powerful method when classification tasks are considered in FL, which has also led to its popularity as the SoTA method in the literature [9, 13]. Nevertheless, we next demonstrate that this method loses its effectiveness when FL is considered for more complex computer vision tasks.
II-B4 FL for Non-Classification Vision Tasks under Class-based Data Heterogeneity
We next take one of the well-known computer vision datasets, called Taskonomy [15], due to the large number of tasks it includes which is why it has been previously used in both centralized [23] and federated [16] multi-task learning as a sole baseline. We utilize the above-described Dirichlet distribution to create a heterogeneous allocation of data across $25$ clients’ local datasets in FL using the available class labels, referred to as “class_scene” in this dataset. [^4] The classes are provided in the dataset as probabilities for each datapoint belonging to each available class from which we choose the class with the highest probability as the determined class for each datapoint. More details on the data processing is provided in Appendix A. In our experiments, we select a representative set of vision tasks in this dataset, which are visualized in Fig. 2 and include “2D Edges,” “Reshading,” “Surface Normals,” “Semantic Segmentation,” “3D Keypoints,” “Euclidean Depth Estimation,” and “Scene Classification.”
These vision tasks – except Scene Classification – can be construed as non-classification tasks since they are associated with a class label. Following the non-iid data induced by Dirichlet distribution across the clients, we study the performance of the FL global model under different levels of heterogeneity – various $\alpha$ values – for the aforementioned vision tasks. The results of this experiment are shown in Fig. 3 (the left plot in each box, titled “Class-based”).
2D Edges — FedAvg Refer to captionRefer to caption
Reshading — FedAvg Refer to captionRefer to caption
Surface Normals — FedAvg Refer to captionRefer to caption
Semantic Segmentation — FedAvg Refer to captionRefer to caption
Euclidean Depth Estimation — FedAvg Refer to captionRefer to caption
3D Keypoints — FedAvg Refer to captionRefer to caption
Scene Classification — FedAvg Refer to captionRefer to caption
Fig. 3: Class-based vs. Embedding-based distribution: Comparison of how performing Dirichlet distribution over the datapoints’ labels (equivalently, the scene class feature) and the extracted embeddings’ affect the performance in FL. As it is shown, distributing the datapoints based on the class/label of the scene they correspond to (the left plot in each box) has not resulted in a significant change in the performance across various values of $\alpha$. However, distributing the datapoints based on the clusters formed by the extracted embeddings (the right plot in each box) with the Dirichlet distributions of various $\alpha$ parameters has created a performance gap across the possible scenarios with the $\alpha=0.1$ (most heterogeneous case) resulting in the worst performance and $\alpha=1000$ (least heterogeneous case) resulting in the best performance. The results further unveil that the class-based data heterogeneity leads to an overestimation of the performance of FL with the loss obtained being smaller than those of embedding-based data heterogeneity. Only the class-based experiments on the Scene Classification task results in a change in performance with increasing $\alpha$. This observation shows that the most distinguishable feature in the Scene Classification task which are the class labels, themselves, are also reflected in the embeddings as the loss decreases are concurrent regarding the loss values for each $\alpha$.
Focusing on the only classification task, i.e., Scene Classification (bottom box), we can see that decreasing the value of $\alpha$ has a notable impact on the performance of the task. However, when focusing on all the other non-classification vision tasks, we can see no significant degeneration in the performance of the global model as the value of $\alpha$ decreases. This key observation signals the ineffectiveness of the class-based (equivalently, label-based) data heterogeneity emulation methods for generic computer vision tasks and calls for revamping the approaches to induce data heterogeneity for such tasks in the FL setting.
Intuitively, class-based data heterogeneity emulation methods often fall short for generic computer vision tasks because non-classification vision tasks perceive data distribution differently from classification tasks. In fact, their perception of data heterogeneity can be label-independent and vary significantly across tasks – an aspect that has largely been overlooked in existing FL literature. For example, consider the Euclidean Depth Estimation task, which focuses on estimating the depth of objects in an image. In this context, two data points with the same label (e.g., both labeled “Living Room”) could exhibit very different depth structures due to variations in camera angle or lighting conditions, making them heterogeneous from the perspective of the Euclidean Depth Estimation task. On the other hand, two data points with different labels (e.g., “Hallway” and “Living Room”) may share similar depth structures, making them homogeneous for this task despite their differing labels. This underscores the need for FL methodologies that account for data heterogeneity from the unique perspectives of each vision task.
Motivated by this observation, we propose an approach to emulate data heterogeneity in FL that accommodates both classification and non-classification computer vision tasks.
Fig. 4: In our approach, data heterogeneity is induced from the unique task perspective. At first, data points are fed to a pre-trained neural network trained on the task (this can be any off-the-shelf model as long as the model is trained for the task of interest). Afterwards, the embeddings of data points are extracted from the penultimate layer of the model. The embeddings are then clustered using clustering algorithms such as K-means, which unveils the similarity/dissimilarity of the data points from the task perspective. Dirichlet distribution is then applied on the clustered data to emulate the local datasets of clients, treating the datapoints in the same cluster as having the same group (analogous to labels in classification tasks).
III Emulation of Data Heterogeneity from the Task Perspective
To address the need for a task-specific modeling of data heterogeneity, we shifted our perspective from the raw input domain to the task output domain in order to get more task-related information. In particular, instead of using the raw data features, such as its label, we aim to use the processed data features, such as embeddings, to emulate data heterogeneity in FL. This stems from the unavailability of representative characteristics, such as class labels, in the case of non-classification tasks, motivating us to identify a utility analogous to class labels that can effectively associate or disassociate datapoints based on the unique perspective of a given task.
In particular, we note that in the case of (deep) neural networks, training the network on different tasks leads to a reconfiguration of its model parameters, where the parameters optimized for each task are often significantly different and unique to that task. Consequently, feeding a raw data point into a neural network trained for a specific task and observing the outputs of different layers can provide unique representations of how the task perceives the data point. Specifically, the outputs of the initial layers, especially in case of convolutional neural networks (CNNs), often correspond to low-level features or simple patterns, while the outputs of the later layers correspond to high-level, task-specific features or other abstract representations. Following this notion, we aimed to extract the task’s perspective of each datapoint by passing each datapoint through a network trained on the task and extracting the output of the penultimate layer as shown in Fig. 4. We will refer to this extracted output as embedding which we use to describe each datapoint from a task’s perspective for the remainder of this discussion.
III-A Task-Specific Data/Embedding Clustering
In a nutshell, in our approach (see Fig. 4), data heterogeneity is induced from the unique perspective of the task through pre-processing the data points. First, data points are fed into a pre-trained neural network trained specifically for the task of interest (any off-the-shelf pre-trained model such as the ResNet[24] or CLIP [25] trained on the target task can be utilized). Next, embeddings for each data point are extracted from the penultimate layer of the model. These embeddings, which encapsulate task-specific features, are then clustered using algorithms such as K-means [26, 27] to reveal the similarity or dissimilarity of data points from the task’s perspective. Finally, a Dirichlet distribution is applied to the clustered data to emulate the local datasets of clients, treating datapoints within the same cluster as belonging to the same group (analogous to “labels” in classification tasks).
Our method enables a task-driven and systematic approach to modeling data heterogeneity in FL, which works both for non-classification as well as classification tasks. In particular, in case of having a classification task, datapoints with the same label are expected to end up in the same cluster/group after conducting the above procedure, which makes class-based data heterogeneity, explained in Sec. II-B, a special case of our method. We demonstrate this via visualizing our method on the Scene Classification task of Taskonomy dataset – which is a classification task – in Fig. 5. As can be seen, the clustering of the data points with respect to their embeddings has resulted in the datapoints being clustered into separable clusters which possess only datapoints from one label. This mapping between the clustering of data points based on their embeddings and their class labels for classification tasks further affirms that the clusters made from the embeddings can model the data points similarity/dissimilarity accurately from the task perspective and the clusters/groups obtained based on the embeddings can be hypothetically considered as having the same labels when applying the Dirichlet distribution.
Fig. 5: The figure illustrates embedding clusters extracted from a Scene Classification task model, where each color represents a specific class. It is evident that the embeddings create a separable space corresponding to the scene classification labels. This demonstrates that distributing data points across FL clients using the clusters/groups depicted in the figure is equivalent to distributing data based on the original scene class labels. The observed alignment between the embedding-based clustering and the class labels underscores the effectiveness of using embeddings as a method to induce task-specific data heterogeneity in FL.
III-B Emulating Heterogeneous Local Datasets for Clients
Formally, once the clustering of datapoints based on their embeddings are obtained, we emulate the local datasets $\mathcal{D}_{n}$ for each client $n$, where the set of ratios for classes $\mathbf{p}$ detailed in Sec. II-B3 is now replaced by the distribution of the embeddings in each cluster/group, which we denote by $\widehat{\mathbf{p}}$. Also, we let the set $\mathcal{G}$ contain the distinct cluster/group indices (e.g., all $16$ groups in Fig. 5). In our approach, each client $n$ will be assigned data from each cluster/group based on its own cluster/group distribution vector $\widehat{\mathbf{q}}_{n}=\mathsf{Dir}(\alpha\widehat{\mathbf{p}})=\{\widehat{q}_{n,g}\}_{g\in\mathcal{G}}$ where $\widehat{q}_{n,g}$ denotes the ratio of client $n$’s dataset containing data from group $g$ (e.g., one of the $16$ groups in Fig. 5). These local datasets are then used during the local training phase of FL.
To demonstrate this revamped notion of data heterogeneity, we revisit the experiments in Sec. II-B to evaluate the effect of the induced data heterogeneity, replacing the class-based method with our embedding-based approach, which groups data points based on their embeddings. The results are presented in Fig. 3 (the right plot in each box, titled “Embedding-based”) next to the class-based results for a better comparison. Unlike the findings from Sec. II-B, where data was distributed across clients based solely on class labels, we observe a pronounced impact of $\alpha$ (which controls the level of data heterogeneity across clients) on the performance of the trained FL global model. Specifically, as data heterogeneity increases (i.e., as $\alpha$ decreases), the global loss converges to a higher value, indicating a degradation in the model’s performance.
These results highlight that task-specific heterogeneity is more effective in modeling a heterogeneous federated system for vision tasks compared to class-based heterogeneity, which has been predominantly used in prior FL studies [10, 5, 6, 11]. Furthermore, the results reveal that the performance of FL for non-classification generic computer vision tasks has been overestimated in previous works [16, 17, 18, 19] due to the reliance on class-based heterogeneity, which fails to adequately emulate the local datasets of clients. Moreover, this experiment underscores that the actual performance and robustness of FL, and its variants, in handling data heterogeneity for generic computer vision tasks remain largely unknown – a critical gap in the literature. We address this gap through the benchmarks and experiments presented in Sec. IV.
In addition to the above-described main takeaways of the results, our approach has broader implications, which we elaborate on through the following remarks.
Remark 1 (Implications on Designing Effective Countermeasures to Address Data Heterogeneity in FL for Vision Tasks).
Our work demonstrates that the prior notion of data heterogeneity, primarily based on the ratio of data points belonging to specific classes, is largely applicable only to classification tasks. By contrast, our method introduces a task-specific approach to inducing data heterogeneity, enabling experiments in FL to be extended to non-classification vision tasks. This shift not only broadens the scope of FL research to include a wider variety of vision tasks but also lays the groundwork for designing vision-specific methods to effectively handle and address/combat data heterogeneity in FL.
Remark 2.
[Implications on the Concept of Task Similarity] Following the footsteps of prior work on modeling task similarity in the centralized ML domain [28, 29], our approach opens a model-driven perspective to quantify the similarity between two tasks. Specifically, this involves clustering data points based on their embeddings obtained from the models pre-trained for each task and comparing the resulting clusters to assess similarity. Using this straightforward approach, we establish a baseline performance for task similarity analysis, with results discussed in Sec. IV-B2. While this serves as a preliminary exploration, more sophisticated methods for task similarity measurement in the FL setting remain an open area for future investigation. Furthermore, this analysis provides an avenue for developing new task grouping strategies in FL, akin to prior efforts in centralized ML [30]. For example, by employing the task similarity measures based on the embedding-based clustering of data points, tasks with highly similar clustering patterns can be grouped and trained collaboratively, potentially enhancing FL performance for related tasks. We leave further exploration of this notion as future work.
Remark 3.
[Implications on Multi-Task FL and its Extensions] A fundamental implications of our method lies in its application to multi-task FL and its extensions. In conventional multi-task FL [19, 18], each client typically has a single local task, with tasks varying across clients. On the other hand, the extended variations of multi-task FL [16, 17] consider scenarios where multiple tasks coexist on a single FL client. In the context of conventional multi-task FL, our method enables researchers to explore how varying levels of data heterogeneity, from the perspectives of different tasks, affect the performance of multi-task FL systems. This opens the door to designing countermeasures specifically aimed at addressing task-induced data heterogeneity in multi-task FL. Similarly, for the extensions of multi-task FL, our method allows researchers to investigate how inducing data heterogeneity from the perspective of one task impacts the performance of other tasks trained concurrently on the same client. Moreover, it facilitates the study of grouping and training local tasks together under different levels of data heterogeneity, potentially leading to new strategies for optimizing task interactions in FL. We leave further exploration of these concepts to future work.
IV Experiments and Benchmarks
In this section, we experiment with various FL methods to give a performance benchmark for each method trained on each vision task under various levels of embedding-based heterogeneity. Our purpose in doing so is three-fold. (i) Universality of the Phenomenon across FL Methods: We demonstrate that the phenomenon observed in Sec. III-B is universally applied across various FL methods and does not depend on the specific FL method employed. (ii) Consistency of Performance Impact on Vision Tasks: We show the performance drop is consistent across all non-classification vision tasks. (iii) Benchmark for Future Research: We provide a benchmark of reference for future endeavors to tackle the task-specific data heterogeneity across non-classification vision tasks in FL.
$\star$ All of our implementations and results are contained in the online GitHub repository of this paper available at https://github.com/KasraBorazjani/task-perspective-het.git
IV-A FL Methods
To observe the effect of embedding-based heterogeneity across various FL methods, we have considered the following six variations of FL:
  1. 1.
    FedAvg[2]: One of the most widely used methods in FL with the dynamics explained in Sec. II-A. The initial results presented in Fig. 3 were also gathered by employing this method.
  2. 2.
    FedProx[3]: FedProx extends FedAvg by incorporating a proximal term into the local objective function to address data heterogeneity across clients. This term penalizes each client’s local model divergence from the global model, encouraging consistency and stability during aggregation.
  3. 3.
    SCAFFOLD[31]: To address heterogeneous data distributions, SCAFFOLD maintains control variates, both on the server and clients, to adjust the local training updates and align them more closely with the global optimization direction. Clients send their updated control variates along with model parameters to the server. The server averages these control variates and redistributes them to clients in subsequent rounds.
  4. 4.
    FedRep[32]: FedRep decouples the optimization of shared and personalized components. Clients optimize a shared representation (encoder) collaboratively while maintaining a personalized classifier layer for task-specific outputs. During interaction with the server, only the shared representation is aggregated, leaving the personalized components untouched. This decoupling is to minimize the negative impact of data heterogeneity while allowing personalization for each client.
  5. 5.
    FedAmp[33]: FedAmp employs a model aggregation strategy that incorporates weighted averaging based on client-specific similarity measures. Clients send their local model updates and similarity scores to the server. The server aggregates these updates, prioritizing contributions from clients with higher similarity to the global objective. The method introduces an adaptive aggregation mechanism that accounts for the variability in clients’ data distributions.
IV-B Further Results and Discussions
To provide insight into our proposed modeling of data heterogeneity, we have conducted two sets of experiments, focusing on different implications of inducing heterogeneity through task perspective, which are
  1. (Set 1)
    Studying the performance of various benchmarks under different levels of data heterogeneity
  2. (Set 2)
    Studying the task similarity through comparing the embedding clusters
These sets of experiments are discussed below in order.
IV-B1 Studying the Performance of Various Benchmarks under Different Levels of Heterogeneity
In order to provide a benchmark of FL performance, we compare the training loss of the benchmarks explained above under class-based and embedding-based data distribution — the same system of $25$ clients as in Sec. II-B4 is considered. This benchmarking provides a baseline performance for each of these SoTA methods under various levels of data heterogeneity. Note that we have shown the results for FedAvg in Fig. 3, which were used to explain our methodology. Subsequently, in the following, we discuss the results for FedProx (see Fig. 6), SCAFFOLD (see Fig. 7), FedRep (see Fig. 8), and FedAmp (see Fig. 9).
Observing the results in Figs. 6-9, we can observe the same phenomenon as in Fig. 3 (see Sec. III-B). In particular, the results indicate that, for the non-classification tasks (i.e., all the tasks except “Scene Classification” presented in the bottom box of each figure), the change in the value of $\alpha$ in the dirichlet distribution in all methods results in a performance gap when our embedding-based distribution of data is utilized (see the right plot in each box). However, at the same time, it can be observed that such a performance gap does not hold for class-based distribution of data (see the left plot in each box). Further, our results verify that embedding-based data heterogeneity and class-based data heterogeneity result in a similar FL performance for the classification task (see the bottom box in Figs. 6-9). The main two takeaways of the results are as follows. (i) These observations further validate that our proposed method of inducing heterogeneity from the tasks’ perspective is capable of creating a heterogeneous data setting more effectively for non-classification vision tasks. (ii) In addition to revealing the performance of all the benchmarks under embedding-based data heterogeneity, the results reveal that all these SoTA methods are subject to overestimation of FL performance under class-based data heterogeneity for vision tasks: when class-based data heterogeneity is induced, the loss is often lower than that of the embedding-based data heterogeneity scenario. In other words, embedding-based data heterogeneity can induce a more pronounced impact on the FL performance (i.e., a higher value of loss).
Furthermore, a major takeaway of implementing these technique were as follows: the tasks exhibit model bias at different numbers of local SGD iterations ($K$), which depends on the task complexity and the underlying data distributions. In other words, some task may need to be trained over longer duration of local training (i.e., higher values of $K$) so as to exhibit the model bias and subsequently a gap in performance under various levels of data heterogeneity. Nevertheless, to have fair comparisons between class-based and embedding-based methods, we maintained the same number of local SGDs per round for both class-based and embedding-based experiments across all benchmark methods and all tasks. The details of these hyperparameter are available at our aforementioned GitHub repository.
IV-B2 Studying the Task Similarity Through Comparing the Embedding Clusters
Revisiting Remarks 2 and 3, it can be noted that the extracted embeddings of the datapoints and the way that they get clustered in our proposed approach may have further use cases. As a preliminary investigation in this direction, we measure the similarities between the embedding clusters, which provide nuanced insights. To this end, we calculate the similarity of embedding clusters across pairs of tasks through the following similarity metric:
$\zeta(t_{1},t_{2};\mathcal{D})=\frac{\sum_{d\in\mathcal{D}}\sum_{d^{\prime}\in\mathcal{D}/\{d\}}\mathbbm{1}_{d,d^{\prime}}(t_{1},t_{2})}{2\binom{|\mathcal{D}|}{2}},$ (4)
where $\mathcal{D}$ denotes the global dataset and $\mathbbm{1}_{d,d^{\prime}}(t_{1},t_{2})$ is a binary indicator that takes the value of one if the embeddings of two datapoints $d$ and $d^{\prime}$ are clustered together or in different clusters under $t_{1}$ and $t_{2}$ tasks. In other words, $\mathbbm{1}_{d,d^{\prime}}(t_{1},t_{2})$ only takes the value of zero when the extracted embeddings of datapoints $d$ and $d^{\prime}$ are clustered together in one of the tasks (i.e., either $t_{1}$ or $t_{2}$) and belong to separate clusters in the other task. The above calculated similarity index will then be a real number in the range $[0,1]$. In essence, higher values of (4) indicate a higher similarity between the clustering configuration of the embeddings extracted for tasks $t_{1}$ and $t_{2}$. [^5] Note that $t_{1}$ and $t_{2}$ will be two of the seven tasks that we considered in this work, summarized in Fig. 2. To provide a holistic view of the similarity of tasks, we measure $\zeta(t_{1},t_{2};\mathcal{D})$ for all the pairs of tasks and then plot the heatmap of these similarities in Table I. In essence, this heatmap reveals how close the perspective of two tasks are towards the datapoints.
Inspecting the results in Table I, the 2D Edges and the Surface Normals tasks have the highest similarity ratio ($0.109$) among the pairs included in the table. This can be intuitionally explained through the similarity between what the two tasks are accomplishing. Specifically, in the 2D Edges task, we aim to find the edges of the objects inside the image which can alternatively be used to discover the separate surfaces inside the image and their normal vectors. On the other hand, the two lowest similarity ratios belong to the 3D Keypoints task when compared with Scene Classification ($0.018$) and Semantic Segmentation tasks ($0.019$). This can also be explained through the fact that 3D Keypoints is a task that considers the very details of the image which is contrary to the higher-level inspection conducted in the Scene Classification and Semantic Segmentation tasks. The rest of the pairwise comparisons fall in between these two extremes. In particular, many of the numerical values are hard to interpret intuitively since human-level visual explanations may not suffice to describe the innate characteristics of these tasks. This makes the reach of the methodology used to obtain the heatmap in Table I more pronouncedly felt in scenarios that human-level intuitions are non-existent. The extensions of the implications of these results to task grouping and multi-task training is left as future work (see Remarks 2 and 3).
TABLE I: Heatmap representation of task relationships/similarities based on the metric $\zeta(t_{1},t_{2};\mathcal{D})$ defined in (4).
Task 1 ($t_{1}$) Task 2 ($t_{2}$)     Scene Classification Surface Normals 2D Edges Euclidean Depth Estimation Semantic Segmentation Reshading 3D Keypoints
Scene Classification     0.0417 0.051 0.048 0.059 0.021 0.0184
Surface Normals     0.041 0.109 0.082 0.045 0.049 0.042
2D Edges     0.051 0.109 0.063 0.064 0.045 0.053
Euclidean Depth Estimation     0.048 0.082 0.063 0.023 0.020 0.035
Semantic Segmentation     0.059 0.045 0.064 0.023 0.028 0.019
Reshading     0.021 0.049 0.045 0.020 0.028 0.054
3D Keypoints     0.018 0.042 0.053 0.035 0.019 0.054
2D Edges — FedProx Refer to captionRefer to caption
Reshading — FedProx Refer to captionRefer to caption
Surface Normals — FedProx Refer to captionRefer to caption
Semantic Segmentation — FedProx Refer to captionRefer to caption
Euclidean Depth Estimation — FedProx Refer to captionRefer to caption
3D Keypoints — FedProx Refer to captionRefer to caption
Scene Classification — FedProx Refer to captionRefer to caption
Fig. 6: Class-based vs. Embedding-based distribution: Comparison of how performing Dirichlet distribution over the datapoints’ labels (equivalently, the scene class feature) and the extracted embeddings’ affect the performance in FL when FedProx methodology is used. The same phenomena observed in Fig. 3 can be observed in the FedProx results.
2D Edges — SCAFFOLD Refer to captionRefer to caption
Reshading — SCAFFOLD Refer to captionRefer to caption
Surface Normals — SCAFFOLD Refer to captionRefer to caption
Semantic Segmentation — SCAFFOLD Refer to captionRefer to caption
Euclidean Depth Estimation — SCAFFOLD Refer to captionRefer to caption
3D Keypoints — SCAFFOLD Refer to captionRefer to caption
Scene Classification — SCAFFOLD Refer to captionRefer to caption
Fig. 7: Class-based vs. Embedding-based distribution: Comparison of how performing Dirichlet distribution over the datapoints’ labels (equivalently, the scene class feature) and the extracted embeddings’ affect the performance in FL when SCAFFOLD methodology is used. The same phenomena observed in Fig. 3 can be observed in the SCAFFOLD results.
2D Edges — FedRep Refer to captionRefer to caption
Reshading — FedRep Refer to captionRefer to caption
Surface Normals — FedRep Refer to captionRefer to caption
Semantic Segmentation — FedRep Refer to captionRefer to caption
Euclidean Depth Estimation — FedRep Refer to captionRefer to caption
3D Keypoints — FedRep Refer to captionRefer to caption
Scene Classification — FedRep Refer to captionRefer to caption
Fig. 8: Class-based vs. Embedding-based distribution: Comparison of how performing Dirichlet distribution over the datapoints’ labels (equivalently, the scene class feature) and the extracted embeddings’ affect the performance in FL when FedRep methodology is used. The same phenomena observed in Fig. 3 can be observed in the FedRep results.
2D Edges — FedAmp Refer to captionRefer to caption
Reshading — FedAmp Refer to captionRefer to caption
Surface Normals — FedAmp Refer to captionRefer to caption
Semantic Segmentation — FedAmp Refer to captionRefer to caption
Euclidean Depth Estimation — FedAmp Refer to captionRefer to caption
3D Keypoints — FedAmp Refer to captionRefer to caption
Scene Classification — FedAmp Refer to captionRefer to caption
Fig. 9: Class-based vs. Embedding-based distribution: Comparison of how performing Dirichlet distribution over the datapoints’ labels (equivalently, the scene class feature) and the extracted embeddings’ affect the performance in FL when FedAmp methodology is used. The same phenomena observed in Fig. 3 can be observed in the FedAmp results.
V Conclusion
In this paper, we have demonstrated that inducing data heterogeneity in FL via a label/class-based approach alone fails to fully encapsulate the real-world data heterogeneity encountered in computer vision tasks beyond classification. We reveal that existing approaches often overestimate FL performance by relying solely on this notion, thereby exposing a critical gap in the literature. To address this limitation, we leveraged pre-trained deep neural networks to extract task-specific data embeddings and proposed a new perspective on data heterogeneity in FL — embedding-based data heterogeneity. By clustering data points based on their embeddings and distributing them among clients using the Dirichlet distribution, we introduce a more nuanced and realistic formulation of data heterogeneity in FL, particularly for non-classification vision tasks. Our experiments validated the impact of this revised framework, providing new benchmark performance measures and offering various nuanced insights. Beyond these contributions, we highlighted several open research directions, paving the way for future advancements in FL under this revised notion of data heterogeneity.
Appendix A Data Preprocessing Details
The data gathered in the Taskonomy dataset pertaining to the scene classification label is provided in the form of a probability vector $\bm{\beta}_{d}=\{\beta_{d,1},\beta_{d,2},\cdots,\beta_{d,J}\}$ for each datapoint $d$. Each $\beta_{d,j},1\leq j\leq J$ represents the probability of the datapoint $d$ belonging to the scene class $j$ in a set of $J=365$ available classes for the datapoints in the Taskonomy dataset. Given the large number of the datapoints available in the dataset, these probabilities were acquired via passing the datapoints through a pre-trained classification model. To simplify the task of distributing the data, we first take the $\max$ across the elements of each $\bm{\beta}_{d}$ for each datapoint and assign it as the class label to that datapoint. However, not all of the $J$ classes have enough datapoints to be distributed across the simulated clients in our experiments. Some classes have small numbers of datapoints which may fall as low as $1$ datapoint in the subset we take from the dataset. Distributing the datapoints belonging to these ill-represented classes across the clients and including them in the classification task will lead to only some clients — as low as one — having data from these classes. This phenomenon will keep us from ever having a homogeneous data distribution across the clients which is one of the settings we have experimented with throughout the paper. Another problem is that training the global model in a federated fashion when one class is very poor in datapoint count will lead to an imbalanced training. Therefore, the biased classifying ability towards the ill-represented class will automatically worsen the results, rendering them unreliable for comparison across different hyperparameter values such as the Dirichlet distribution’s concentration parameter $\alpha$. To overcome this obstacle, we take the top 16 most-populated classes into consideration in our experiments and distribute their data across the clients. We use the same set of data for all tasks in order to ensure no effect from the less populated classes is induced into the results of the experiments in any of the tasks.